# ConFit v2: Improving Resume-Job Matching using Hypothetical Resume Embedding and Runner-Up Hard-Negative Mining

### v1

@@ 1 Introduction
Since both resume and job posts are often stored as text data, there has been an increased interest in using transformer models to model resume-job fit (or referred to as “person-job fit”). Many prior works Zhu et al. (2018); Qin et al. (2018); Bian et al. (2020); Yang et al. (2022); Shao et al. (2023) focus on designing complex modeling techniques to model resume-job matching. For example, APJFNN Qin et al. (2018) uses hierarchical recurrent neural networks to process the job and resume content, and DPGNN Yang et al. (2022) uses a dual-perspective graph neural network to model the relationship between resumes and jobs. Unlike these manual model/feature engineering methods, more recently Anonymous (2024) shows that dense retrieval techniques like contrastive learning achieve significant improvements without relying on any complex designs. By introducing dense retrieval methods to resume-job matching, ConFit Anonymous (2024) presented a simple yet effective method to rank thousands of resumes/jobs in milliseconds, by combining training neural encoders such as E5 Wang et al. (2022) with inner-product search algorithms such as FAISS Johnson et al. (2019).
@@ 2 Background
A resume-job matching (or often called person-job fit) system models the compatibility between a resume $R$ and a job post $J$, allowing the systems to recommend the most suitable candidates for a given job, or suggest the most relevant jobs for a given candidate Bian et al. (2020); Yang et al. (2022); Shao et al. (2023); Anonymous (2024). Since resumes and job posts are often stored as text data, many prior works consider using neural networks (e.g., encoders) to quantify the compatibility between resumes and jobs:
@@
where $f_{\theta}$ could be directly modeled by a neural network Zhu et al. (2018); Yang et al. (2022); Shao et al. (2023), or by computing inner product/cosine similarity Anonymous (2024) between a resume embedding $f_{\theta}(R)$ and a job embedding $f_{\theta}(J)$.
Despite the rapid growth of online recruitment platforms and the increasing availability of job posts and resume data, there are very few interaction labels between any resumes and job posts Bian et al. (2020); Anonymous (2024). This is because a candidate usually applies to only a few positions, and a job interviewer often reviews only a few resumes for a given job post. Often, the resulting dataset $\mathcal{D}=\{R_{i},J_{i},y_{i}\}$ has size $|\mathcal{D}|\ll n_{R}\times n_{J}$, where $n_{R}$ and $n_{J}$ are the total number of resumes and jobs respectively, and $y_{i}\in\{0,1\}$ is a binary signal representing whether a resume $R_{i}$ is accepted for an interview by a job $J_{i}$. For example, in both the IntelliPro and AliYun datasets (Section 4.1) used in this work, less than 0.05% of the total possible (resume, job) pairs are annotated. This label sparsity poses challenges in: 1) crafting high-quality training data/hard negatives for training a neural encoder $f_{\theta}$; and 2) learning a representation space that generalizes well across diverse resumes and job posts.
@@ 3 Approach
We propose ConFit v2, a simple and general-purpose approach that improves ConFit Anonymous (2024) to model resume-job compatibility. Similar to ConFit, we use an encoder to produce an embedding of a given resume or job post, and model the matching score between an $\langle R,J\rangle$ pair as the cosine similarity of their representations. Unlike ConFit, ConFit v2 uses a simpler encoder architecture, and substantially improves ranking performance by using 1) hypothetical reference resume generated by an LLM to help improve the encoder’s job data understanding; and 2) a novel hard-negative mining method to enhance encoder contrastive learning. We provide a high-level overview of ConFit v2 during inference in Figure 2. We detail each modification below.
@@ 3.1 Encoder Architecture
Many prior works on resume-job matching employ complex neural architectures to encode human-designed matching heuristics Zhu et al. (2018); Yang et al. (2022). For example, Shao et al. (2023); Anonymous (2024) uses a hierarchical attention mechanism to model interactions between each section (e.g., Education, Experience, etc.) of a resume and a job post. However, in practice, these modifications often rely on specific domain knowledge and require additional structural constraints on resume/job data, restricting their generalizability.
@@ 3.3 Runner-Up Mining
To accurately model resume-job compatibility, the encoder needs to distinguish between matching and near-matching resume-job pairs. ConFit Anonymous (2024) achieves state-of-the-art performance by training the encoder with contrastive loss, where given a job the model uses 1) random resumes as in-batch negatives and 2) rejected resumes as hard negatives, and vice versa for resumes. However, due to label scarcity (Section 2), the quantity of rejected resumes/hard negatives is highly limited.
@@ 3.4 ConFit v2
To train ConFit v2 encoder, we use HyRe to augment all job posts with generated reference resumes (Section 3.2), and replace them with original job posts for later training and testing. Then, we use RUM to mine hard-negative resumes and (augmented) jobs (Section 3.3). Finally, we train the simplified ConFit using the modified contrastive learning loss $\mathcal{L}$ from Anonymous (2024):
@@ 4.2 Implementation Details
We follow ConFit Anonymous (2024) and experiment with two encoder architectures. We consider E5-base Wang et al. (2022), and Jina-v2-base. Mohr et al. (2024). Both encoders are trained on large-scale Chinese-English bilingual text data, and were amongst the best embedding models according to benchmarks such as MTEB Muennighoff et al. (2022) at the time of the project. For HyRe, we use GPT-4o-mini as the LLM for generating hypothetical resumes due to its cost-effectiveness. We provide the prompts used in Section B.1. We then train ConFit with HyRe using contrastive learning (without hard negatives), and use to mine hard negatives with RUM. For RUM, we use the top 3%-4% percentile to pick hard-negatvies across all experiments. Finally, we train ConFit again, using 1) HyRe augmented job data; and 2) RUM mined hard negatives. We use Eq (1) with a batch size of 4 and 2 hard negatives per batch, and a learning rate of 1e-5. This is the final encoder used for ConFit v2. We run all experiments using one A100 80GB GPU. To speed up cosine similarity computation across a large pool of resumes and jobs, we use FAISS Johnson et al. (2019) throughout all experiments.
@@ 4.3 Baselines
Prior person-job fit systems include MV-CoN Bian et al. (2020), InEXIT Shao et al. (2023), and ConFit Anonymous (2024) and more Qin et al. (2018); Zhu et al. (2018); Yang et al. (2022). MV-CoN considers a co-teaching network Han et al. (2018) to learn from sparse, noisy person-job fit data, and InEXIT uses hierarchical attention to model interactions between the text fields of a resume-job pair. ConFit focuses on dense retrieval techniques, and uses contrastive learning and LLM-based data augmentation to train an encoder. Other older methods such as APJFNN Qin et al. (2018), PJFNN Zhu et al. (2018), and DPGNN Yang et al. (2022) are omitted since they are already outperformed by these more recent methods.
@@ 4.5 Iterative RUM
We present the results in Figure 3. In general, we find that training with more than one iteration ($t>1$) improves resume ranking performance compared to a single iteration ($t=1$). However, as the number of iterations increases, improvement begins to drop. We believe this is similar to many model self-improvement research Huang et al. (2023); Yu et al. (2024). While model-created data often helps improve performance, these data also contains noises that can compound over multiple iterations. We believe techniques such as model ensembles Zhang et al. (2011); de Souza P. Moreira et al. (2024) could help mitigate this issue, and we leave this for future work.
@@ 4.6 Ablation Studies
Table 3 presents our ablation studies for each component of ConFit v2. First, we consider only simplifying the encoder architecture (denoted as ConFit), and use rejected resumes/jobs as hard negatives used by ConFit Anonymous (2024). We find that this simplification significantly improves ranking performance. We believe this is due to the strong performance of recent embedding models, which can effectively learn complex interactions between resumes and job posts directly from data. Next, we use HyRe to generate reference resumes, augment job posts, and train ConFit on the augmented data (denoted as +HyRe). We find this to improve 2.9% absolute in recall and 3.1% nDCG scores on average. Finally, we use RUM to mine hard resume/(augmented) job negatives (denoted as +RUM), and replace them with the rejected resume/jobs used by ConFit. In Table 3, we find that RUM further improved by, on average, 1.2% absolute in recall scores, and 1.8% absolute in nDCG across both datasets. We find that all components are important for ConFit v2.
@@ 5.1 Error Analysis
To analyze the errors made by ConFit v2, we manually inspect 30 negative resume-job pairs from the ranking tasks that are incorrectly ranked at top 5% and are before at least one positive pair. For each incorrectly ranked pair, we compare against other positive resume-job pairs from the dataset, and categorize the errors following the criteria from Anonymous (2024): unsuitable, where some requirements in the job post are not satisfied by the resume; less competent, where a resume satisfies all job requirements, but many competing candidates have a higher degree/more experience; out-of-scope, where a resume satisfies all requirements, appears competitive compared to other candidates, but is still rejected due to other (e.g., subjective) reasons not presented in our resume/job data themselves; and potentially suitable, where a resume from the ranking tasks satisfied the requirements and seemed competent, but had no label in the original dataset.
@@ 6 Related Work
Early person-job fit systems that use neural networks typically focus on architecture designs. These methods include Qin et al. (2018); Zhu et al. (2018); Rezaeipourfarsangi and Milios (2023); Jiang et al. (2020); Mhatre et al. (2023), which explores architectures such as RNN, LSTM Staudemeyer and Morris (2019) and CNN O’Shea and Nash (2015). However, these are significantly outperformed by more recent transformer-based methods such as Maheshwary and Misra (2018); Rezaeipourfarsangi and Milios (2023); Bian et al. (2019); Zhang et al. (2023), which focuses on small architecture or loss modifications. For example, MV-CoN Bian et al. (2020) uses a co-teaching network Malach and Shalev-Shwartz (2018) to perform gradient updates based on model’s confidence to data noises; InEXIT Shao et al. (2023), which uses hierarchical attention to model resume-job interactions; and DPGNN Yang et al. (2022), which uses a graph-based approach with a novel BPR loss to optimize for resume/job ranking. However, these methods tend to focus on task-specific modifications, such as assuming no unseen resumes/jobs at test time or assuming access to internal data such as whether the recruiter sent private messages to the job applicant. More recently, ConFit Anonymous (2024) focuses on applying dense retrieval techniques to person-job fit. Using contrastive learning with encoders such as E5 Wang et al. (2022), Anonymous (2024) demonstrates its performance and flexibility by achieving the best scores in almost all ranking and classification tasks across two different person-job fit datasets. In this work, we propose HyRe and RUM to further enhance ConFit v2, achieving an average absolute improvement of 13.8% in recall and 17.5% in nDCG across job ranking and resume-ranking tasks in two resume-job benchmarks.
@@
To further improve retrieval results, researchers considered methods such as query/document expansions Carpineto and Romano (2012) and hard-negative mining Robinson et al. (2021). Recent query/document expansion approaches include HyDE Gao et al. (2022) and HyQE Zhou et al. (2024), which prompts an LLM to augment the original query or the passage to retrieve and is typically used without finetuning (the LLM or the encoder model). Common hard-negative mining strategies often use BM25 to select top-k unlabeled/incorrect samples as hard negatives Karpukhin et al. (2020); Zhao et al. (2024); Nguyen et al. (2023). Other methods include: ANCE Xiong et al. (2021) which samples negatives from the top retrieved documents asynchronously during training; SimANS Zhou et al. (2022) which samples hard negatives using a manually designed probability distribution; and more Zhan et al. (2021); Li et al. (2024). ConFit Anonymous (2024) presents the first successful attempt to use contrastive learning for person-job fit. We extend ConFit and adapt recent dense retrieval techniques to person-job fit by 1) sampling hard-negative resumes/jobs from a “runner-up” percentile ranges to mitigate selecting false-negatives; and 2) few-shot prompting an LLM to augment a job posted used for later encoder training.
@@ 8 Limitations
To our knowledge, there is no standardized, public person-job fit dataset [^5] The AliYun dataset used in this work is no longer publicly available as of 09-11-2023. that can be used to compare performances of existing systems Zhu et al. (2018); Qin et al. (2018); Bian et al. (2020); Yang et al. (2022); Shao et al. (2023). This is due to the highly sensitive nature of resume and job post content, making large-scale person-job fit datasets largely proprietary. We follow Anonymous (2024) and provide the best effort to make ConFit v2 reproducible and extensible for future work. We will open-source implementations of ConFit v2, related baselines, data processing scripts, and dummy train/valid/test data files that can be used to test drive our system end-to-end. We will also privately release our model weights and full datasets to researchers under appropriate license agreements, aiming to make future research in person-job fit more accessible.
@@ 9 Ethical Considerations
ConFit v2 uses and trains pretrained encoders such as E5 Wang et al. (2022) and Jina-v2 Günther et al. (2023), and it is well-known that these models contain biases including but not limited to gender, race, demographics, and more Brunet et al. (2019); May et al. (2019); Jentzsch and Turan (2022); Caliskan et al. (2022). For practical person-job fit systems, we believe it is crucial to ensure that they do not exhibit these biases, such as preferring a certain gender for certain jobs. In this work, we have followed best practices from prior work Yang et al. (2022); Anonymous (2024) to remove sensitive information (e.g., gender) from our data. However, our bias analysis (Section 5.2) reveals that gender disparities in the trained model could stem from imbalances in the dataset. We do not condone using ConFit v2 for real-world applications without using de-biasing techniques such as Bolukbasi et al. (2016); Cheng et al. (2021); Gaci et al. (2022); Guo et al. (2022); Schick et al. (2021), and in general, we do not condone the use of ConFit v2 for any morally unjust purposes. To our knowledge, there is little work on investigating or mitigating biases in existing person-job fit systems, and we believe this is a crucial direction for future person-job research.

### v2

@@ 1 Introduction
Since both resume and job posts are often stored as text data, there has been an increased interest in using transformer models to model resume-job fit (or referred to as “person-job fit”). Many prior works Zhu et al. (2018); Qin et al. (2018); Bian et al. (2020); Yang et al. (2022); Shao et al. (2023) focus on designing complex modeling techniques to model resume-job matching. For example, APJFNN Qin et al. (2018) uses hierarchical recurrent neural networks to process the job and resume content, and DPGNN Yang et al. (2022) uses a dual-perspective graph neural network to model the relationship between resumes and jobs. Unlike these manual model/feature engineering methods, more recently Yu et al. (2024b) shows that dense retrieval techniques like contrastive learning achieve significant improvements without relying on any complex designs. By introducing dense retrieval methods to resume-job matching, ConFit Yu et al. (2024b) presented a simple yet effective method to rank thousands of resumes/jobs in milliseconds, by combining training neural encoders such as E5 Wang et al. (2022) with inner-product search algorithms such as FAISS Johnson et al. (2019).
@@ 2 Background
A resume-job matching (or often called person-job fit) system models the compatibility between a resume $R$ and a job post $J$, allowing the systems to recommend the most suitable candidates for a given job, or suggest the most relevant jobs for a given candidate Bian et al. (2020); Yang et al. (2022); Shao et al. (2023); Yu et al. (2024b). Since resumes and job posts are often stored as text data, many prior works consider using neural networks (e.g., encoders) to quantify the compatibility between resumes and jobs:
@@
where $f_{\theta}$ could be directly modeled by a neural network Zhu et al. (2018); Yang et al. (2022); Shao et al. (2023), or by computing inner product/cosine similarity Yu et al. (2024b) between a resume embedding $f_{\theta}(R)$ and a job embedding $f_{\theta}(J)$.
Despite the rapid growth of online recruitment platforms and the increasing availability of job posts and resume data, there are very few interaction labels between any resumes and job posts Bian et al. (2020); Yu et al. (2024b). This is because a candidate usually applies to only a few positions, and a job interviewer often reviews only a few resumes for a given job post. Often, the resulting dataset $\mathcal{D}=\{R_{i},J_{i},y_{i}\}$ has size $|\mathcal{D}|\ll n_{R}\times n_{J}$, where $n_{R}$ and $n_{J}$ are the total number of resumes and jobs respectively, and $y_{i}\in\{0,1\}$ is a binary signal representing whether a resume $R_{i}$ is accepted for an interview by a job $J_{i}$. For example, in both the IntelliPro and AliYun datasets (Section 4.1) used in this work, less than 0.05% of the total possible (resume, job) pairs are annotated. This label sparsity poses challenges in: 1) crafting high-quality training data/hard negatives for training a neural encoder $f_{\theta}$; and 2) learning a representation space that generalizes well across diverse resumes and job posts.
@@ 3 Approach
We propose ConFit v2, a simple and general-purpose approach that improves ConFit Yu et al. (2024b) to model resume-job compatibility. Similar to ConFit, we use an encoder to produce an embedding of a given resume or job post, and model the matching score between an $\langle R,J\rangle$ pair as the cosine similarity of their representations. Unlike ConFit, ConFit v2 uses a simpler encoder architecture, and substantially improves ranking performance by using 1) hypothetical reference resume generated by an LLM to help improve the encoder’s job data understanding; and 2) a novel hard-negative mining method to enhance encoder contrastive learning. We provide a high-level overview of ConFit v2 during inference in Figure 2. We detail each modification below.
@@ 3.1 Encoder Architecture
Many prior works on resume-job matching employ complex neural architectures to encode human-designed matching heuristics Zhu et al. (2018); Yang et al. (2022). For example, Shao et al. (2023); Yu et al. (2024b) uses a hierarchical attention mechanism to model interactions between each section (e.g., Education, Experience, etc.) of a resume and a job post. However, in practice, these modifications often rely on specific domain knowledge and require additional structural constraints on resume/job data, restricting their generalizability.
@@ 3.3 Runner-Up Mining
To accurately model resume-job compatibility, the encoder needs to distinguish between matching and near-matching resume-job pairs. ConFit Yu et al. (2024b) achieves state-of-the-art performance by training the encoder with contrastive loss, where given a job the model uses 1) random resumes as in-batch negatives and 2) rejected resumes as hard negatives, and vice versa for resumes. However, due to label scarcity (Section 2), the quantity of rejected resumes/hard negatives is highly limited.
@@ 3.4 ConFit v2
To train ConFit v2 encoder, we use HyRe to augment all job posts with generated reference resumes (Section 3.2), and replace them with original job posts for later training and testing. Then, we use RUM to mine hard-negative resumes and (augmented) jobs (Section 3.3). Finally, we train the simplified ConFit using the modified contrastive learning loss $\mathcal{L}$ from Yu et al. (2024b):
@@ 4.2 Implementation Details
We follow ConFit Yu et al. (2024b) and experiment with two encoder architectures. We consider E5-base Wang et al. (2022), and Jina-v2-base. Mohr et al. (2024). Both encoders are trained on large-scale Chinese-English bilingual text data, and were amongst the best embedding models according to benchmarks such as MTEB Muennighoff et al. (2022) at the time of the project. For HyRe, we use GPT-4o-mini as the LLM for generating hypothetical resumes due to its cost-effectiveness. We provide the prompts used in Section B.1. We then train ConFit with HyRe using contrastive learning (without hard negatives), and use to mine hard negatives with RUM. For RUM, we use the top 3%-4% percentile to pick hard-negatvies across all experiments. Finally, we train ConFit again, using 1) HyRe augmented job data; and 2) RUM mined hard negatives. We use Eq (1) with a batch size of 4 and 2 hard negatives per batch, and a learning rate of 1e-5. This is the final encoder used for ConFit v2. We run all experiments using one A100 80GB GPU. To speed up cosine similarity computation across a large pool of resumes and jobs, we use FAISS Johnson et al. (2019) throughout all experiments.
@@ 4.3 Baselines
Prior person-job fit systems include MV-CoN Bian et al. (2020), InEXIT Shao et al. (2023), and ConFit Yu et al. (2024b) and more Qin et al. (2018); Zhu et al. (2018); Yang et al. (2022). MV-CoN considers a co-teaching network Han et al. (2018) to learn from sparse, noisy person-job fit data, and InEXIT uses hierarchical attention to model interactions between the text fields of a resume-job pair. ConFit focuses on dense retrieval techniques, and uses contrastive learning and LLM-based data augmentation to train an encoder. Other older methods such as APJFNN Qin et al. (2018), PJFNN Zhu et al. (2018), and DPGNN Yang et al. (2022) are omitted since they are already outperformed by these more recent methods.
@@ 4.5 Iterative RUM
We present the results in Figure 3. In general, we find that training with more than one iteration ($t>1$) improves resume ranking performance compared to a single iteration ($t=1$). However, as the number of iterations increases, improvement begins to drop. We believe this is similar to many model self-improvement research Huang et al. (2023); Yu et al. (2024a). While model-created data often helps improve performance, these data also contains noises that can compound over multiple iterations. We believe techniques such as model ensembles Zhang et al. (2011); de Souza P. Moreira et al. (2024) could help mitigate this issue, and we leave this for future work.
@@ 4.6 Ablation Studies
Table 3 presents our ablation studies for each component of ConFit v2. First, we consider only simplifying the encoder architecture (denoted as ConFit), and use rejected resumes/jobs as hard negatives used by ConFit Yu et al. (2024b). We find that this simplification significantly improves ranking performance. We believe this is due to the strong performance of recent embedding models, which can effectively learn complex interactions between resumes and job posts directly from data. Next, we use HyRe to generate reference resumes, augment job posts, and train ConFit on the augmented data (denoted as +HyRe). We find this to improve 2.9% absolute in recall and 3.1% nDCG scores on average. Finally, we use RUM to mine hard resume/(augmented) job negatives (denoted as +RUM), and replace them with the rejected resume/jobs used by ConFit. In Table 3, we find that RUM further improved by, on average, 1.2% absolute in recall scores, and 1.8% absolute in nDCG across both datasets. We find that all components are important for ConFit v2.
@@ 5.1 Error Analysis
To analyze the errors made by ConFit v2, we manually inspect 30 negative resume-job pairs from the ranking tasks that are incorrectly ranked at top 5% and are before at least one positive pair. For each incorrectly ranked pair, we compare against other positive resume-job pairs from the dataset, and categorize the errors following the criteria from Yu et al. (2024b): unsuitable, where some requirements in the job post are not satisfied by the resume; less competent, where a resume satisfies all job requirements, but many competing candidates have a higher degree/more experience; out-of-scope, where a resume satisfies all requirements, appears competitive compared to other candidates, but is still rejected due to other (e.g., subjective) reasons not presented in our resume/job data themselves; and potentially suitable, where a resume from the ranking tasks satisfied the requirements and seemed competent, but had no label in the original dataset.
@@ 6 Related Work
Early person-job fit systems that use neural networks typically focus on architecture designs. These methods include Qin et al. (2018); Zhu et al. (2018); Rezaeipourfarsangi and Milios (2023); Jiang et al. (2020); Mhatre et al. (2023), which explores architectures such as RNN, LSTM Staudemeyer and Morris (2019) and CNN O’Shea and Nash (2015). However, these are significantly outperformed by more recent transformer-based methods such as Maheshwary and Misra (2018); Rezaeipourfarsangi and Milios (2023); Bian et al. (2019); Zhang et al. (2023), which focuses on small architecture or loss modifications. For example, MV-CoN Bian et al. (2020) uses a co-teaching network Malach and Shalev-Shwartz (2018) to perform gradient updates based on model’s confidence to data noises; InEXIT Shao et al. (2023), which uses hierarchical attention to model resume-job interactions; and DPGNN Yang et al. (2022), which uses a graph-based approach with a novel BPR loss to optimize for resume/job ranking. However, these methods tend to focus on task-specific modifications, such as assuming no unseen resumes/jobs at test time or assuming access to internal data such as whether the recruiter sent private messages to the job applicant. More recently, ConFit Yu et al. (2024b) focuses on applying dense retrieval techniques to person-job fit. Using contrastive learning with encoders such as E5 Wang et al. (2022), Yu et al. (2024b) demonstrates its performance and flexibility by achieving the best scores in almost all ranking and classification tasks across two different person-job fit datasets. In this work, we propose HyRe and RUM to further enhance ConFit v2, achieving an average absolute improvement of 13.8% in recall and 17.5% in nDCG across job ranking and resume-ranking tasks in two resume-job benchmarks.
@@
To further improve retrieval results, researchers considered methods such as query/document expansions Carpineto and Romano (2012) and hard-negative mining Robinson et al. (2021). Recent query/document expansion approaches include HyDE Gao et al. (2022) and HyQE Zhou et al. (2024), which prompts an LLM to augment the original query or the passage to retrieve and is typically used without finetuning (the LLM or the encoder model). Common hard-negative mining strategies often use BM25 to select top-k unlabeled/incorrect samples as hard negatives Karpukhin et al. (2020); Zhao et al. (2024); Nguyen et al. (2023). Other methods include: ANCE Xiong et al. (2021) which samples negatives from the top retrieved documents asynchronously during training; SimANS Zhou et al. (2022) which samples hard negatives using a manually designed probability distribution; and more Zhan et al. (2021); Li et al. (2024). ConFit Yu et al. (2024b) presents the first successful attempt to use contrastive learning for person-job fit. We extend ConFit and adapt recent dense retrieval techniques to person-job fit by 1) sampling hard-negative resumes/jobs from a “runner-up” percentile ranges to mitigate selecting false-negatives; and 2) few-shot prompting an LLM to augment a job posted used for later encoder training.
@@ 8 Limitations
To our knowledge, there is no standardized, public person-job fit dataset [^5] The AliYun dataset used in this work is no longer publicly available as of 09-11-2023. that can be used to compare performances of existing systems Zhu et al. (2018); Qin et al. (2018); Bian et al. (2020); Yang et al. (2022); Shao et al. (2023). This is due to the highly sensitive nature of resume and job post content, making large-scale person-job fit datasets largely proprietary. We follow Yu et al. (2024b) and provide the best effort to make ConFit v2 reproducible and extensible for future work. We will open-source implementations of ConFit v2, related baselines, data processing scripts, and dummy train/valid/test data files that can be used to test drive our system end-to-end. We will also privately release our model weights and full datasets to researchers under appropriate license agreements, aiming to make future research in person-job fit more accessible.
@@ 9 Ethical Considerations
ConFit v2 uses and trains pretrained encoders such as E5 Wang et al. (2022) and Jina-v2 Günther et al. (2023), and it is well-known that these models contain biases including but not limited to gender, race, demographics, and more Brunet et al. (2019); May et al. (2019); Jentzsch and Turan (2022); Caliskan et al. (2022). For practical person-job fit systems, we believe it is crucial to ensure that they do not exhibit these biases, such as preferring a certain gender for certain jobs. In this work, we have followed best practices from prior work Yang et al. (2022); Yu et al. (2024b) to remove sensitive information (e.g., gender) from our data. However, our bias analysis (Section 5.2) reveals that gender disparities in the trained model could stem from imbalances in the dataset. We do not condone using ConFit v2 for real-world applications without using de-biasing techniques such as Bolukbasi et al. (2016); Cheng et al. (2021); Gaci et al. (2022); Guo et al. (2022); Schick et al. (2021), and in general, we do not condone the use of ConFit v2 for any morally unjust purposes. To our knowledge, there is little work on investigating or mitigating biases in existing person-job fit systems, and we believe this is a crucial direction for future person-job research.